OpenStudy (anonymous):

SOLVE: 10x + 1+ 9x -11=180

OpenStudy (anonymous):

x=10

OpenStudy (anonymous):

thank you (: can u show the steps of u did it though?

OpenStudy (anonymous):

o,man,its hard to put what I did in words,cause I did it in my head lol

OpenStudy (anonymous):

well i know that 10x and 9x is 19x but what about the 1 and 11? thats what im stuck on

OpenStudy (anonymous):

x=190/19 =10

OpenStudy (anonymous):

\[10x +1 +9x-11 = 180\] \[19x-10=180\] \[19x=180+10\] \[19x=190\] \[x=10\]

OpenStudy (anonymous):

Oh! thanks for that^ (:

OpenStudy (anonymous):

If you`re talking to me you're welcome.

OpenStudy (anonymous):

yeah i was (: